OpenStudy (anonymous):

is this right? Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (5, 5) m = -3 y=-3x+20

OpenStudy (anonymous):

Again !: Y2-y=m(X2-x) 5-y=-3(5-x) 5-y=-15+3x -y=-20+3x y=20-3x And right .

OpenStudy (anonymous):

this is how i did.... y-5=-3(x-5) y-5=-3x+15 add 5 0n both sides y=-3x+20????????????????????